### v1

@@ 5 Conclusion
Our evaluation shows that ViTs outperform state-of-the-art CNN baselines—achieving up to a 4% increase in overall accuracy and similar gains across other key metrics. Beyond raw accuracy, ViTs generalise with a broader habitat scene understanding, demonstrating superior interpretability over CNNs. We also find that SupCon is a more effective training paradigm than classical supervised learning for training the model to distinguish between visually similar habitats. SupCon considerably reduces misclassifications between commonly confused habitat pairs—such as Improved and Neutral Grassland—outperforming standard supervised learning. Finally, our best ViT model not only exceeds CNN performance but also matches photographic identification by experienced ecological experts on key classification metrics, highlighting the potential of our work to reduce the need for human experts to interpret habitats at this level in the field.
@@
Declaration of Generative AI and AI-assisted Technologies in the Writing Process
@@ Appendix A Embedding Analysis
We employ UMAP, CHI, and DBI to analyse the embedding space produced by the trained image encoder, both qualitatively and quantitatively. UMAP is applied to compress the embeddings extracted from the encoder and visualise them in a 2D graph. Embeddings from different classes can be clustered on the UMAP, showing whether they are heavily overlapped or distinctively separated, patterns that can impact the classification performance. UMAP is particularly useful for us to observing the effect of the chosen learning methods, the conventional supervised learning and supervised contrastive learning. In addition, we further utilise the Calinski–Harabasz (CH) index (Caliński and Harabasz, 1974), and Davies–Bouldin (DB) index (Davies and Bouldin, 1979) to quantitatively evaluate the inter-class variance and intra-class variance in the embedding space. CH index and DB index are both popular metrics for evaluating clustering performance, rewarding methods that are capable of achieving more separable clusters for features. A higher CH index or a lower DBI implies a better clustering quality, with low intra-class variances and high inter-class variances.
@@
Figure 8: UMAPs of the embedding space produced by the image encoder on test set. SupCon is found to generate tightly clusters that are more separable from each other compared to supervised learning. Habitat symbols in this figure: $\bullet$ Acid Grassland;$\bullet$ Arable and Horticulture;$\bullet$ Bog; $\bullet$ Bracken; $\bullet$ Broadleaved Mixed and Yew Woodland; $\bullet$ Calcareous Grassland;$\bullet$ Coniferous Woodland; $\bullet$ Dwarf Shrub Heath; $\bullet$ Fen, Marsh, Swamp; $\bullet$ Improved Grassland; $\bullet$ Inland Rock; $\bullet$ Littoral Sediment; $\bullet$ Montane; $\bullet$ Neutral Grassland; $\bullet$ Supra-littoral Sediment; $\bullet$ Urban.
@@
[…]
@@
Appendix B Analysis of the Human-Expert Benchmark
@@
Figures 9, 10, 11, 12 present confusion matrices for three human experts and our best model (SwinT-B + SupCon) on the representative CS subset described in Section 3.4.3 of the main paper, allowing a habitat-by-habitat comparison at Level 3. Four main observations are provided:
@@
Appendix C Habitat Description in UKHab
Table 6: Habitat descriptions defined in UKHab (UKHab Ltd, 2023). Habitats in bold belong to level 2, followed by their level 3 subcategories.
@@ Appendix C Habitat Description in UKHab
Acid Grassland (L3) Vegetation dominated by grasses and herbs on a range of lime-deficient soils that have been derived from acidic bedrock or from superficial deposits such as sands and gravels. Such soils usually have a low base-status, with a pH of $<5.5$.
Bracken (L4) [^1] Bracken is treated as a L3 habitat in the CS dataset as explained in Section 3.1 Land with Bracken Pteridium aquilinum at $>95\%$ canopy cover at the height of the growing season.
@@
Improved Grassland (L3)  [^2] Improved Grassland is a Broad Habitat that originates from the UK BAP Priority Habitats system (Joint Nature Conservation Committee, n.d.) and is used as equivalent to an L3 habitat (Modified Grassland) in UKHab. This broad habitat type is characterised by vegetation dominated by a few fast-growing grasses on fertile, neutral soils. It is frequently characterised by an abundance of rye-grass Lolium spp. and white clover Trifolium repens. Improved grasslands are typically either managed as pasture or mown regularly for silage production or in non-agricultural contexts for recreation and amenity purposes.
[…]
@@
Broadleaved and Mixed Woodland (L3) Vegetation dominated by trees that are $>5\,\text{m}$ high when mature, which form a distinct, although sometimes open, canopy with a canopy cover of $>25\%$. It includes stands of both native and non-native broadleaved tree species and Yew Taxus baccata where the percentage cover of these trees in the stand is $>20\%$ of the total cover of the trees present.
Coniferous Woodland (L3) Vegetation dominated by trees that are $>5\,\text{m}$ high when mature, which form a distinct, although sometimes open, canopy that has a cover of $>25\%$. It includes stands of both native and non-native coniferous tree species (with the exception of Yew Taxus baccata) where the percentage cover of these trees in the stand is $>80\%$ of the total cover of the trees present.
Heathland and Shrub (L2) Vegetation with a $>25\%$ cover of dwarf shrub species that are $<1.5\,\text{m}$ high or woody species $\leq 5\,\text{m}$ high.
Dwarf Shrub Heath (L3) Vegetation that has a $>25\%$ cover of plant species from the heath family (ericoids), Dwarf Gorse Ulex minor or Western Gorse Ulex gallii.
@@
Sparsely Vegetated Land (L2) Unvegetated, disturbed (regularly or drastically periodically) or sparsely vegetated habitats (permanently or periodically naturally unvegetated areas) that are inhabited by stress-tolerating vegetation with cover $<50\%$.
@@
Montane (L3) [^4] Montane is a Broad Habitat that originates from the UK BAP Priority Habitats system (Joint Nature Conservation Committee, n.d.) and is used as equivalent to an L3 habitat in UKHab. We refer to Natural England (2014) for its definition. Montane habitats consist of a range of near-natural vegetation which lie above the natural treeline. The vegetation within these habitats includes dwarf-shrub heaths, grass-heaths, dwarf-herb communities, willow scrub, and snowbed communities. The most abundant vegetation types are heaths dominated by heather Calluna vulgaris and billbery Vaccinium myrtillus, typically with abundant bryophytes and/or lichens; and siliceous alpine and boreal grasslands with stiff sedge Carex bigelowii and moss heaths. Rarer vegetation types include snow-bed communities with dwarf willow Salix herbacea and various bryophytes and lichens, and sub-arctic willow scrub

### v2

@@ 5 Conclusion
Our evaluation shows that ViTs outperform state-of-the-art CNN baselines—achieving up to a 4% increase in overall accuracy and similar gains across other key metrics. Beyond raw accuracy, ViTs generalise with a broader habitat scene understanding, demonstrating superior interpretability over CNNs. We also find that SupCon is a more effective training paradigm than classical supervised learning for training ViT to distinguish between visually similar habitats. SupCon considerably reduces misclassifications between commonly confused habitat pairs—such as Improved and Neutral Grassland—outperforming standard supervised learning. Finally, our best ViT model not only exceeds CNN performance but also matches photographic identification by experienced ecological experts on key classification metrics, highlighting the potential of our work to reduce the need for human experts to interpret habitats at this level in the field.
@@
Funding
This work was funded by Natural Environment Research Council of UK Research and Innovation [grant numbers UKRI054].
Declaration of Generative AI and AI‐assisted Technologies in the Writing Process
@@
Data Availability
The code used for model training and evaluation is publicly available at https://github.com/WhiteGiveFive/Habitat-Classification-from-Ground-Level-Imagery (MIT license). The data used in this study are not publicly hosted due to UKCEH’s data privacy policy but may be obtained from UKCEH under a research licence upon request. Requests for access to the ground-level photographs and corresponding UKHab labels should be directed to Dr Lisa Norton (lrn@ceh.ac.uk).
@@ Appendix A Embedding Analysis
We employ UMAP, CHI, and DBI to analyse the embedding space produced by the trained image encoder, both qualitatively and quantitatively. UMAP is applied to compress the embeddings extracted from the encoder and visualise them in a 2D graph. Embeddings from different classes can be clustered on the UMAP, showing whether they are heavily overlapped or distinctively separated, patterns that can impact the classification performance. UMAP is particularly useful for us to observing the effect of the chosen learning methods, the conventional supervised learning and supervised contrastive learning. In addition, we further utilise the Calinski–Harabasz (CH) index (Caliński01011974), and Davies–Bouldin (DB) index (4766909) to quantitatively evaluate the inter-class variance and intra-class variance in the embedding space. CH index and DB index are both popular metrics for evaluating clustering performance, rewarding methods that are capable of achieving more separable clusters for features. A higher CH index or a lower DBI implies a better clustering quality, with low intra-class variances and high inter-class variances.
@@
[…]
@@
CH Index $\uparrow\uparrow$ DB Index $\downarrow\downarrow$
@@
Appendix B Effects of SupCon on CNN Backbones
For our employed CNNs, SupCon reduces misclassification for certain classes but increases it for others, leading to minimal or even negative overall improvement. Notably, the confusion matrices (Figures 2, 3, 4) show that SupCon boosts recall for Neutral Grassland but increases false positives, especially making more Improved Grassland misclassified as Neutral. In our case, SupCon tightens within-class embeddings but also pulls some samples toward incorrect clusters.
Our underlying hypotheses for this observation are twofold. (i) Due to locality bias, CNNs may discard visual cues needed to discriminate habitats, yielding embeddings that are indistinguishable. (ii) Some anchor samples are poor class representatives for contrastive learning (e.g., images that plausibly resemble both Neutral and Improved Grassland). To address (i), we can replace CNNs with ViTs to obtain globally informed, more expressive embeddings that better suit SupCon. To address (ii), we can actively select anchors with low model uncertainty (e.g., minimum-entropy criterion) for SupCon.
Figure 2: Delta CM of WRN-50-2 with SupCon: based on the CM produced by SupCon.
Figure 3: Delta CM of ResNeXt50 with SupCon: based on the CM produced by SupCon.
Figure 4: Delta CM of EfficientNetV2-L with SupCon: based on the CM produced by SupCon.
Appendix C Analysis of the Human-Expert Benchmark
@@
Figures 5, 6, 7, 8 present confusion matrices for three human experts and our best model (SwinT-B + SupCon) on the representative CS subset described in Section 3.4.3 of the main paper, allowing a habitat-by-habitat comparison at Level 3. Four main observations are provided:
@@
Appendix D Habitat Description in UKHab
Table 2: Habitat descriptions defined in UKHab (UKHab2023). Habitats in bold belong to level 2, followed by their level 3 subcategories.
@@ Appendix D Habitat Description in UKHab
Acid Grassland (L3) Vegetation dominated by grasses and herbs on a range of lime-deficient soils that have been derived from acidic bedrock or from superficial deposits such as sands and gravels. Such soils usually have a low base-status, with a pH of <5.5<$5.$
Bracken (L4) [^1] Bracken is treated as a L3 habitat in the CS dataset as explained in Section 3.1 Land with Bracken Pteridium aquilinum at $95\%$ canopy cover at the height of the growing season.
@@
Improved Grassland (L3)  [^2] Improved Grassland is a Broad Habitat that originates from the UK BAP Priority Habitats system (jncc_bap_habitats) and is used as equivalent to an L3 habitat (Modified Grassland) in UKHab. This broad habitat type is characterised by vegetation dominated by a few fast-growing grasses on fertile, neutral soils. It is frequently characterised by an abundance of rye-grass Lolium spp. and white clover Trifolium repens. Improved grasslands are typically either managed as pasture or mown regularly for silage production or in non-agricultural contexts for recreation and amenity purposes.
Neutral Grassland (L3) Vegetation dominated by grasses and herbs on a range of neutral soils, usually with a pH of $4.5\text{--}6.5$.
@@
Broadleaved and Mixed Woodland (L3) Vegetation dominated by trees that are $5\,\text{m}$ high when mature, which form a distinct, although sometimes open, canopy with a canopy cover of $25\%$. It includes stands of both native and non-native broadleaved tree species and Yew Taxus baccata where the percentage cover of these trees in the stand is $20\%$ of the total cover of the trees present.
Coniferous Woodland (L3) Vegetation dominated by trees that are $5\,\text{m}$ high when mature, which form a distinct, although sometimes open, canopy that has a cover of $25\%$. It includes stands of both native and non-native coniferous tree species (with the exception of Yew Taxus baccata) where the percentage cover of these trees in the stand is $80\%$ of the total cover of the trees present.
Heathland and Shrub (L2) Vegetation with a $25\%$ cover of dwarf shrub species that are $1.5\,\text{m}$ high or woody species $\leq 5\,\text{m}$ high.
Dwarf Shrub Heath (L3) Vegetation that has a $25\%$ cover of plant species from the heath family (ericoids), Dwarf Gorse Ulex minor or Western Gorse Ulex gallii.
@@
Sparsely Vegetated Land (L2) Unvegetated, disturbed (regularly or drastically periodically) or sparsely vegetated habitats (permanently or periodically naturally unvegetated areas) that are inhabited by stress-tolerating vegetation with cover $50\%$.
@@
Montane (L3) [^4] Montane is a Broad Habitat that originates from the UK BAP Priority Habitats system (jncc_bap_habitats) and is used as equivalent to an L3 habitat in UKHab. We refer to NaturalEngland2014Montane for its definition. Montane habitats consist of a range of near-natural vegetation which lie above the natural treeline. The vegetation within these habitats includes dwarf-shrub heaths, grass-heaths, dwarf-herb communities, willow scrub, and snowbed communities. The most abundant vegetation types are heaths dominated by heather Calluna vulgaris and billbery Vaccinium myrtillus, typically with abundant bryophytes and/or lichens; and siliceous alpine and boreal grasslands with stiff sedge Carex bigelowii and moss heaths. Rarer vegetation types include snow-bed communities with dwarf willow Salix herbacea and various bryophytes and lichens, and sub-arctic willow scrub